OpenStudy (anonymous):

Use the given information to find the exact value of the expression. sin α = 21/29, α lies in quadrant II, and cos β = 15/17 , β lies in quadrant I Find sin (α - β)

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

draw a triangle, label the sides

OpenStudy (anonymous):



OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

then you need \(\cos(\alpha)\) which you find via pythagoras

OpenStudy (anonymous):

I got 475/493

OpenStudy (anonymous):

you mean as a final answer?

OpenStudy (anonymous):

i didn't get anywhere near there yet i am still at \[\cos(\alpha)=-\frac{20}{29}\]

OpenStudy (anonymous):

yes

OpenStudy (mathmale):

Hints: 1) You're going to need the "difference formula," sin (a-b) = sin a cos b -cos a sin b. 2) From the given info, you're going to need to find \[\cos \alpha~and ~\sin \beta \]

OpenStudy (mathmale):

I'd suggest you begin with (2) and then focus on (1).